One number is 3 less than another. Their sum is 15. Find the numbers.
Answer by checkley71(8403)

X+(X+3)=15 OR 2X+3=15 OR 2X=12 OR X=6 THE X+3=9 THEN THE TWO NUMBERS ARE 6 & 9